OpenStudy (anonymous):

sin(37.5)sin(7.5) (1/2)cos(30)-cos(45) then what?

OpenStudy (anonymous):

i dont quite understand ur answers dont match up

OpenStudy (anonymous):

sin(37.5)sin(7.5)=(1/2)cos(30)-cos(45) im using the product to sum identity (1/2)[cos(x-y)-cos(x+y)]

OpenStudy (anonymous):

i c, so u want it simplified?

OpenStudy (anonymous):

\[(1/2) [(\sqrt{3}/2)-(\sqrt{2}/2)]\]

OpenStudy (anonymous):

yes please

jimthompson5910 (jim_thompson5910):

hint: use the identity I've attached

OpenStudy (anonymous):

im using this identity tho

OpenStudy (anonymous):

the answer is (sqrt(3)-sqrt(2))/4

jimthompson5910 (jim_thompson5910):

oh ok that will work too

OpenStudy (anonymous):

remember what cos 30 is equivalent to and cos 45?

OpenStudy (anonymous):

lol looking up u already did it

OpenStudy (anonymous):

why'd u ask if u knew?

OpenStudy (anonymous):

cos(30)=\[\sqrt{3}/2\] cos(45)=\[\sqrt{2}/2\]

OpenStudy (anonymous):

i just didnt know how to simplify it thanks tho

OpenStudy (anonymous):

take out 1/2 out of both of them np